# A compact QUBO encoding of computational logic formulae demonstrated on cryptography constructions

### v1

@@ 3.3 QUBO encoding of CNF, DNF and ANF forms
$r=1$ when the multi-variable $XOR$ clause is satisfied, and $r=0$ otherwise. It is obvious from identities that bidirectional equality can be expressed with XNOR (i.e. NOT XOR) as given by $(r\leftrightarrow SAT)\equiv[(1-r)\oplus SAT]$. Therefore, to include an extra bit $r$ for the result of a multi-input XOR clause one merely encode the extended
@@
SAT expression with the strategy described earlier in Sec. 3.1. When the multi-input XOR defined on the variables $x_{i}$ ($10\leq i\leq N-$ 1) evaluates to true (false), (26) can be only satisfied with $r=1$ ($r=0$).
@@
For odd $N$ one extra (linear) term is required containing the last variable $x_{N}-1$ and the constant in the front should be increased by $1$.
We further prove that introducing substitutions for higher order products in k-DNF one can simply reuse the existing k-CNF method of Ref.[61]. In this context the terminology of ”higher order product” means multi-input conjunctions with at least $3$ inputs. The CNF method in [61] is based on the simple transformation similar to the one used in Tseitin encoding to 3-CNF. In other words, first we assign a new substitution variable to each of the disjunctive terms in the CNF via a strong equivalence. The introduced substitution variables are then used in a conjunctive relation encode-able via (28). For example, having a 3-CNF given by
@@
We interpret Eq. (32) in the following way: on the left side one have a disjunctive clause (or equivalently a multi-input OR function) with $k+1$ variables. On the right side we have a conjunction of $2$-input disjunctive clauses, which is nothing else than a 2-CNF expression. Since the clauses in a 2-CNF formula are directly encode-able via QUBO without introducing further substitution variables, there is only the $(k+1)$-CNF clause to be encoded on the left side of (32). To this end we can use the technique described in Sec. 3.2. For encoding the $2$-input disjunctive clauses on the right side $1+rx_{i}-r-x_{i}$ can be used. To encode the conjunction of these terms for all possible $i$ one can sum up these expressions resulting in:
@@
(We notice that the formula (28) to encode the conjunctive part of this expression does not work, as it would give cubic terms in the encoding.) To satisfy the SAT expression (32), both sides need to evaluate to true. This translates to a unified QUBO encoding as the sum of the QUBO encoding of the left and right sides. The QUBO instance takes on its global minimum of 00, when both components reaches their global minimum at the same time. In summary, to encode a whole k-CNF expression into a QUBO form we make use of the substitution variables introduced by the strong equivalence (30) for the $k$-term sub-clauses and encode the outer conjunctive relation between all of these variables labeled by $r_{p}$. Here, the index $p$ was introduced to label the individual disjunctive terms in the k-CNF. The conjunctive relation between the variables $r_{p}$ is then encoded via Eq. (28).
@@ 4 QUBO Encoding Techniques for Cryptography Functions
Multi-input NAND: ¯$\overline{\bigwedge}X$ Eqs. (19), (23) and (24) with $L=0$ and $H=N-1$
@@ 4.1 Advanced Encryption Standard (AES)
A natural application that arises containing an abundance of XOR constraints is the Advanced Encryption Standard (AES) [68, 69] which beyond the non-linear affine transformations inside of the substitution boxes (S-boxes), consists entirely of XOR operations, or with other words, additions in Galois field with two elements (GF($2$)). This is thematic throughout many symmetric cryptography schemes, in fact.
@@
AES is a cryptography scheme using a symmetric key, that is the same key is used to both decoding and encoding. The key size used for an AES cipher specifies the number of transformation rounds that convert the input into the final encoded output. The number of rounds are $10$ rounds for 128-bit keys, $12$ rounds for 192-bit keys and $14$ rounds for 256-bit keys. The individual rounds break down to $4$ steps performed on an input of $4\;\textrm{byte}\times 4\;\textrm{byte}$ data blocks of the input: (i) substituting each of the bytes via a substitution box (S-box), (ii) shifting rows in the $4\;\textrm{byte}\times 4\;\textrm{byte}$ matrix, (iii) mixing columns in this matrix and finally, (iv) add a cipher key to the resulting elements of the matrix with bit-wise XOR operation. In each round the output of the previous round is processed with a newly generated $16$ byte key ((iv)-th step) obtained from the initial AES key (which is either $128$, $192$ or $256$ bit long) via a key schedule technique. At the beginning, before performing any rounds, an instance of step (iv) is applied on the input. The last round omits step (iii). In order to recover the initial key and decode the message, one might follow the route of translating the AES transformation rounds into a boolean circuit (i.e. a circuit composed from boolean functions like the AND, OR or XOR gates) which can be further processed into a QUBO form using the developed techniques in this work. During this procedure the bits of the key are considered to be unknown, while the bits of the input text and the encrypted data are known constants. The global minimum of the QUBO will satisfy the boolean circuit and the encryption key can be recovered. For AES, the S-box represents the additional challenge for optimal QUBO encoding. This could be achieved by designing an efficient boolean circuit of the multiplication inverse function over GF($2^{8}$) and than encode it into a QUBO form. However, searching for an optimal boolean circuit representation is an NP hard problem in itself [70]. An efficient decomposition of the multiplication inverse over GF($2^{8}$) was given by Canright in [71]. According to these results, the multiplication inverse in GF($2^{8}$) can be expressed via $3$ multiplications in GF($2^{4}$) and a single multiplication inverse in GF($2^{4}$). Additionally, a GF($2^{4}$) multiplication is decomposed in terms of $3$ GF($2^{2}$) multiplications. Finally, the multiplication in GF($2^{2}$) is reduced enough in complexity to be expressed via optimal boolean circuit. The AND gates in the circuit of the multiplication appear minimally three times, defining the multiplicative complexity of this operation to $3$. Three GF($2^{2}$) multiplications also occur in one use of the inverse modulo in GF($2^{4}$) as well. In summary, the boolean circuit of a GF($2^{8}$) has a complexity of $3\times 3\times 3+3\times 3=36$ AND gates. All of the inputs to these gates require XOR substitutions (described in Sec. 3.3), though there are several which are re-usable. However, we note an observation similar to that of Boyar and Peralta [72] that the multiplication inverse of GF($2^{4}$) has an exhaustively provable multiplicative complexity of $5$ AND gates which is a decent savings from the original $3\times 3=9$ AND gates. (The exhaustive search is carried out with the procedure described in Sec. 2 via the polynomial $g$, the AND gates in $g$ are equivalent to quadratic multiplications. We discuss the details of this procedure later in this section.) In fact, to date $3\times 3\times 3+5=32$ is the best known result for the number of non-XOR gates required in the circuit of the GF($2^{8}$) multiplication inverse circuit as reported more recently in [73]. However, it remains unclear if multiplication in GF($2^{4}$) and multiplication inverse in GF($2^{8}$) have optimal multiplicative complexities of $9$ and $32$ if expressed via boolean circuits, or less. This is because determining the multiplicative complexity of boolean functions is considered intractable beyond $6$ variables [70]. In line with this finding, we experienced these problems to be extremely challenging in the context of our exhaustive search technique presented in Sec. 2 and failed to tackle either of them. Any remaining details in QUBO encoding of AES (steps (ii)-(iv) in the individual rounds) involve a reuse and simplification of the XOR formulae developed in Sec. 3.1. In order to showcase the potential in optimizing a set of multi-input XOR formulae we outline the following example. Having two XOR expressions $z_{0}\leftrightarrow(a\oplus b\oplus c\oplus d$ and $z_{1}\leftrightarrow(a\oplus b\oplus c\oplus e)$ we can reuse $z_{0}$ in the second expression as $z_{1}\leftrightarrow(z_{0}\oplus d\oplus e)$ and reduce the expression in input variable count. To encode a shorter multi-input XOR expression we need to cover a smaller range with the substitution variables, hence saving variables in the encoding. As the optimal reduction of these XOR formulae is a parity-style variant of the NP-hard minimum set covering problem, we opted a heuristic using a greedy algorithm to first find the largest reusable XOR substitution, then the next largest one, etc. We also placed substitution variables on all variables at the input to the multiplication of GF($2^{8}$) beyond the substitutions for the involved functions of multiplication and associated multiplication inverse in GF($2^{4}$). It should be noted that on order of hundreds to thousands of XOR counting set substitution variables can be saved by careful heuristics.
@@
Now we discuss in more details the procedure to solve the QUBO encoding problem of the multiplication function in GF($2^{2}$) and multiplication inverse in GF($2^{4}$) needed to encode the S-box. The multiplication function in GF($2^{2}$) calculates the product of two elements in GF($2^{2}$), both described via two bits. Therefore the multiplication maps from two bit-pairs onto a single bit-pair: $(x_{1},x_{0},y_{1},y_{0})\rightarrow(z_{1},z_{0})$. The function is fully represented via the look-up-table between the input-output bits. The $\mathbf{x}$ bit-string (first input in the polynomial $g$ given by Eq. (6)) has $6$ elements. The bit combinations in the look-up-table determine the min-terms for which the QUBO encoding should give the global minimum 00. For any other bit-string combination the polynomial $g$ needs to be greater than 00. To find the suitable coefficients we solve the ILP problem given by Eq. (8). The ILP could be solved with two substitution variables $s_{1}$ and $s_{2}$, giving the following QUBO encoding:
@@
To encode the multiplication inverse of GF($2^{4}$), we follow a similar procedure. In this case we are dealing with a map with a single four-bit input and a four-bit output: $(x_{1},x_{0},x_{2},x_{3})\rightarrow(z_{3},z_{2},z_{1},z_{0})$. Therefore, the bit-string $\mathbf{x}$ (again, $\mathbf{x}$ labels the input of the polynomial $g$) has $8$ elements, for which we could solve the associated ILP problem with a single substitution variable $s$, giving a QUBO form:
@@
Table 3: The number of variables in the QUBO encoding of several Cryptography Functions. The results also indicate the sparseness of the resulting QUBO matrix (the density is calculated as a ratio of the nonzero elements in the QUBO matrix and $N^{2}$, with $N$ being the number of variables), and the magnitude of the largest element in the matrix. The encoding results for MD5, SHA1 and SHA-256 are presented for $2$ cases: (a) using blocks of size $B=6$ for modular addition and (b) $B=1$ with k-SAT/ANF forms partitioned into at most $13$ clauses (to keep the coefficients low), see the main text for explanation. For encoding with $B=6$ and for AES there was no limitation on the length of the SAT forms.
@@ 4.2 Encoding constructions of MD5, SHA-1 and SHA-256 hash functions
As the main result of our work is the QUBO encoding of the AES we omit the detailed mathematical description of these hash algorithms. We just mention that they rely on a modular addition over GF($2^{32}$) as well as $3$ and $4$-input XOR operations, ternary conditional functions (ternary conditional function operates on $3$ inputs: one binary input is used to choose between the two remaining inputs.) and majority conditional functions (MCF), like the one defined on $3$ input variables: $MCF(x,y,z)=(x\wedge y)\vee(y\wedge z)\vee(z\wedge x)$. In other words, the MCF returns $1$ if at least two of the input variables are $1$, and 00 otherwise. The key point in encoding the mentioned hash functions is the efficient encoding of these components. The $4$-input XOR (used in SHA-256) with substitution variable encoding its result (see Eq. (25)) can be encoded with the procedure described in Sec. 3.3. For the $3$-input XOR formula $z\leftrightarrow x_{0}\oplus x_{1}\oplus x_{2}$ we provide an alternative QUBO encoding with a single substitution variable $(x_{0}+x_{1}+x_{2}-z+2s-2)^{2}$ obtained by ILP. The weak substitution variable takes on the value $1$ and minimizes the formula when $x_{0}+x_{1}+x_{2}+1-z=1$. The ternary selection $z\leftrightarrow(x_{0}\wedge x_{1})\vee(\overline{x_{0}}\wedge x_{2})$ is encoded via a QUBO form:
@@
The encoding obtained by solving the associated ILP resulted in a single substitution variable $s$. It takes on the value $1$ when $x_{0}+x_{1}+x_{2}+1-z=1$. Additionally, the $3$-input majority function $z\leftrightarrow x_{0}\wedge x_{1}\vee x_{0}\wedge x_{2}\vee x_{1}\wedge x_{2}$ can be encoded without any substitution variable via the QUBO form
@@
The substitution variable $s$ takes on the value $1$ and minimizes the formula when $\overline{x_{1}}\wedge(x_{0}+x_{2}+x_{3}\overset{?}{=}1)$ is True. Finally, only the discussion to encode the modular addition remains. Generally speaking, the most efficient way to encode addition in a QUBO is to write the addition as an equality $a+b=c$, then squaring the difference of the left and right sides and take the minimum of the quadratic function: $(a+b-c)^{2}\rightarrow 0$. When expanding $a$, $b$ and $c$ into binary formats, we obtain the QUBO representation of the addition. The drawback of this simple approach are the large coefficients in front of the most significant bit-variables. When dealing with $32$-bit addition relevant in the case of the studied hash functions, these coefficients become to large to be applicable on QUBO solving hardware. To address this issue, Ref. [12] developed a block-wise strategy to break the QUBO encoding of additions into smaller blocks. The QUBO formulae of the blocks are inter-connected via carry bits, as the sum of the terms might overflow the range covered by the bits of the block. Following this concept, we split the inputs into blocks by
@@
where $B\geq 1$ ($B\in\mathbb{N}$) is the block-size. With other words, the value of $B$ tells us the number of ”ordinary” bits encompassed via a single coefficient $a_{i}$ ($10\leq i\leq i_{max}-$ 1 with $i_{max}=\lceil\log_{2}(a)\rceil/B$, given that $B$ divides $\lceil\log_{2}(a)\rceil$ with no remainder) in a range $10\leq a_{i}\leq 2^{B}-$ 1. By expanding $b$ and $c$ as well, the addition reads as
@@
At this point, however, we need to account for potential overflows in the additions $a_{i}+b_{i}-c_{i}$ as none of the coefficients is allowed to get larger than $12^{B}-$ 1. This can be accomplished by introducing carry (or substitution) variables labeled by $s_{i}$, encoding the potential overflow of the sub-addition $a_{i}+b_{i}$. With this logic Eq. (46) is equivalent to the formulation
@@
In the MD5 algorithm the modular additions can be grouped into a single four-input and one two-input addition over 32-bit numbers in each round. The addition with $4$ inputs implies the modification of (47) to
$\begin{array}[]{ll}0=a_{0}+b_{0}+c_{0}+d_{0}-e_{0}-&\textrm{if: }i=0\\ -2^{B}s_{0}-2^{B+1}s_{1}&\\ 0=a_{i}+b_{i}+c_{i}+d_{i}-e_{i}+s_{2i-2}+&\textrm{if: }0<i\leq i_{max}-1\\ +2s_{2i-1}-2^{B}s_{2i}-2^{B+1}s_{2i+1}&\end{array}$ (48)
@@
describing the four-term addition $a+b+c+d=e$. The $4$ terms in the addition imply $2$ extra bits in each block to cover the potential overflow in the addition. The lower bound of maximal coefficient in the QUBO encoding due to expanding the squares lead to $2^{2B+2}$. We have also tried to further simplify the encoding logic with the ILP strategy developed in Sec. 2, though did not find any better encoding with less variables. In the actual MD5 implementation some pre-computed constants are also used in the procedure opening the door for possible reductions in the variable count by substitution constant bits and merging QUBO terms. Because of this, the coefficient in front of the merged QUBO terms are expected to further increase by some extent. We summarize our results on MD5 QUBO encoding in Table. 3. The encoding with $B=6$ (MD5a) seems to be more efficient than with $B=1$ (MD5b). This is because the $B=1$ case involves roughly $6$-times more carry bits than the encoding with $B=6$ sized blocks. Also, a limit on the maximal length of ANF forms (i.e. the maximal number of inputs in XOR operations) has been set to $k=26$ for the case of $B=1$. (For the $B=6$ case and for the AES QUBO encodings we left the length of the SAT forms unlimited.) This little adjustment can be also used to decrease the coefficients in the QUBO instance, as the substitution variables would need to cover smaller ranges. This good practice becomes especially valuable when needed to port the QUBO instance onto a machine with limited numerical precision. Longer ANF forms are partitioned into smaller segments via substitution variables according to the procedure in Sec. 3.3. Finally, we need to mention that savings in the variable count usually comes with the price of larger coefficients, as reported in Table. 3. To run calculations on an actual hardware a reasonable trade-off needs to be found between QUBO variable count and the magnitude of the coefficients. Comparing to the QUBO encoding of Ref. [36] having $17280$ QUBO variables, our encoding strategy brings in a significant reduction to $10272$ QUBO variables obtained with $B=6$ sized blocks for addition.

### v2

@@ 3.3 QUBO encoding of CNF, DNF and ANF forms
$r=1$ when the multi-variable XORXOR clause is satisfied, and $r=0$ otherwise. It is obvious from identities that bidirectional equality can be expressed with XNOR (i.e. NOT XOR) as given by $(r\leftrightarrow SAT)\equiv[(1-r)\oplus SAT]$. Therefore, to include an extra bit $r$ for the result of a multi-input XOR clause one merely encode the extended
@@
SAT expression with the strategy described earlier in Sec. 3.1. When the multi-input XOR defined on the variables $x_{i}$ ($0\leq i\leq N-1$) evaluates to true (false), (26) can be only satisfied with $r=1$ ($r=0$).
@@
For odd NN one extra (linear) term is required containing the last variable $x_{N}-1$ and the constant in the front should be increased by 11.
We further prove that introducing substitutions for higher order products in k-DNF one can simply reuse the existing k-CNF method of Ref.[61]. In this context the terminology of ”higher order product” means multi-input conjunctions with at least 33 inputs. The CNF method in [61] is based on the simple transformation similar to the one used in Tseitin encoding to 3-CNF. In other words, first we assign a new substitution variable to each of the disjunctive terms in the CNF via a strong equivalence. The introduced substitution variables are then used in a conjunctive relation encode-able via (28). For example, having a 3-CNF given by
@@
We interpret Eq. (32) in the following way: on the left side one have a disjunctive clause (or equivalently a multi-input OR function) with $k+1$ variables. On the right side we have a conjunction of 22-input disjunctive clauses, which is nothing else than a 2-CNF expression. Since the clauses in a 2-CNF formula are directly encode-able via QUBO without introducing further substitution variables, there is only the $(k+1)$-CNF clause to be encoded on the left side of (32). To this end we can use the technique described in Sec. 3.2. For encoding the 22-input disjunctive clauses on the right side $1+rx_{i}-r-x_{i}$ can be used. To encode the conjunction of these terms for all possible $i$ one can sum up these expressions resulting in:
@@
(We notice that the formula (28) to encode the conjunctive part of this expression does not work, as it would give cubic terms in the encoding.) To satisfy the SAT expression (32), both sides need to evaluate to true. This translates to a unified QUBO encoding as the sum of the QUBO encoding of the left and right sides. The QUBO instance takes on its global minimum of 0, when both components reaches their global minimum at the same time. In summary, to encode a whole k-CNF expression into a QUBO form we make use of the substitution variables introduced by the strong equivalence (30) for the $k$-term sub-clauses and encode the outer conjunctive relation between all of these variables labeled by $r_{p}$. Here, the index $p$ was introduced to label the individual disjunctive terms in the k-CNF. The conjunctive relation between the variables $r_{p}$ is then encoded via Eq. (28).
@@ 4 QUBO Encoding Techniques for Cryptography Functions
Multi-input NAND: $\overline{\bigwedge}X$ Eqs. (19), (23) and (24) with $L=0$ and H=N1H=N-1
@@ 4.1 Advanced Encryption Standard (AES)
A natural application that arises containing an abundance of XOR constraints is the Advanced Encryption Standard (AES) [68, 69] which beyond the non-linear affine transformations inside of the substitution boxes (S-boxes), consists entirely of XOR operations, or with other words, additions in Galois field with two elements (GF(22)). This is thematic throughout many symmetric cryptography schemes, in fact.
@@
AES is a cryptography scheme using a symmetric key, that is the same key is used to both decoding and encoding. The key size used for an AES cipher specifies the number of transformation rounds that convert the input into the final encoded output. The number of rounds are $10$ rounds for 128-bit keys, $12$ rounds for 192-bit keys and $14$ rounds for 256-bit keys. The individual rounds break down to 44 steps performed on an input of $4\;\textrm{byte}\times 4\;\textrm{byte}$ data blocks of the input: (i) substituting each of the bytes via a substitution box (S-box), (ii) shifting rows in the $4\;\textrm{byte}\times 4\;\textrm{byte}$ matrix, (iii) mixing columns in this matrix and finally, (iv) add a cipher key to the resulting elements of the matrix with bit-wise XOR operation. In each round the output of the previous round is processed with a newly generated $16$ byte key ((iv)-th step) obtained from the initial AES key (which is either $128$, $192$ or $256$ bit long) via a key schedule technique. At the beginning, before performing any rounds, an instance of step (iv) is applied on the input. The last round omits step (iii). In order to recover the initial key and decode the message, one might follow the route of translating the AES transformation rounds into a boolean circuit (i.e. a circuit composed from boolean functions like the AND, OR or XOR gates) which can be further processed into a QUBO form using the developed techniques in this work. During this procedure the bits of the key are considered to be unknown, while the bits of the input text and the encrypted data are known constants. The global minimum of the QUBO will satisfy the boolean circuit and the encryption key can be recovered. For AES, the S-box represents the additional challenge for optimal QUBO encoding. This could be achieved by designing an efficient boolean circuit of the multiplication inverse function over GF($2^{8}$) and than encode it into a QUBO form. However, searching for an optimal boolean circuit representation is an NP hard problem in itself [70]. An efficient decomposition of the multiplication inverse over GF($2^{8}$) was given by Canright in [71]. According to these results, the multiplication inverse in GF($2^{8}$) can be expressed via 33 multiplications in GF($2^{4}$) and a single multiplication inverse in GF($2^{4}$). Additionally, a GF($2^{4}$) multiplication is decomposed in terms of 33 GF($2^{2}$) multiplications. Finally, the multiplication in GF($2^{2}$) is reduced enough in complexity to be expressed via optimal boolean circuit. The AND gates in the circuit of the multiplication appear minimally three times, defining the multiplicative complexity of this operation to 33. Three GF($2^{2}$) multiplications also occur in one use of the inverse modulo in GF($2^{4}$) as well. In summary, the boolean circuit of a GF($2^{8}$) has a complexity of $3\times 3\times 3+3\times 3=36$ AND gates. All of the inputs to these gates require XOR substitutions (described in Sec. 3.3), though there are several which are re-usable. However, we note an observation similar to that of Boyar and Peralta [72] that the multiplication inverse of GF($2^{4}$) has an exhaustively provable multiplicative complexity of 55 AND gates which is a decent savings from the original $3\times 3=9$ AND gates. (The exhaustive search is carried out with the procedure described in Sec. 2 via the polynomial $g$, the AND gates in $g$ are equivalent to quadratic multiplications. We discuss the details of this procedure later in this section.) In fact, to date $3\times 3\times 3+5=32$ is the best known result for the number of non-XOR gates required in the circuit of the GF($2^{8}$) multiplication inverse circuit as reported more recently in [73]. However, it remains unclear if multiplication in GF($2^{4}$) and multiplication inverse in GF($2^{8}$) have optimal multiplicative complexities of 99 and $32$ if expressed via boolean circuits, or less. This is because determining the multiplicative complexity of boolean functions is considered intractable beyond 66 variables [70]. In line with this finding, we experienced these problems to be extremely challenging in the context of our exhaustive search technique presented in Sec. 2 and failed to tackle either of them. Any remaining details in QUBO encoding of AES (steps (ii)-(iv) in the individual rounds) involve a reuse and simplification of the XOR formulae developed in Sec. 3.1. In order to showcase the potential in optimizing a set of multi-input XOR formulae we outline the following example. Having two XOR expressions $z_{0}\leftrightarrow(a\oplus b\oplus c\oplus d$ and $z_{1}\leftrightarrow(a\oplus b\oplus c\oplus e)$ we can reuse $z_{0}$ in the second expression as $z_{1}\leftrightarrow(z_{0}\oplus d\oplus e)$ and reduce the expression in input variable count. To encode a shorter multi-input XOR expression we need to cover a smaller range with the substitution variables, hence saving variables in the encoding. As the optimal reduction of these XOR formulae is a parity-style variant of the NP-hard minimum set covering problem, we opted a heuristic using a greedy algorithm to first find the largest reusable XOR substitution, then the next largest one, etc. We also placed substitution variables on all variables at the input to the multiplication of GF($2^{8}$) beyond the substitutions for the involved functions of multiplication and associated multiplication inverse in GF($2^{4}$). It should be noted that on order of hundreds to thousands of XOR counting set substitution variables can be saved by careful heuristics.
@@
Now we discuss in more details the procedure to solve the QUBO encoding problem of the multiplication function in GF($2^{2}$) and multiplication inverse in GF($2^{4}$) needed to encode the S-box. The multiplication function in GF($2^{2}$) calculates the product of two elements in GF($2^{2}$), both described via two bits. Therefore the multiplication maps from two bit-pairs onto a single bit-pair: $(x_{1},x_{0},y_{1},y_{0})\rightarrow(z_{1},z_{0})$. The function is fully represented via the look-up-table between the input-output bits. The $\mathbf{x}$ bit-string (first input in the polynomial $g$ given by Eq. (6)) has 66 elements. The bit combinations in the look-up-table determine the min-terms for which the QUBO encoding should give the global minimum 0. For any other bit-string combination the polynomial $g$ needs to be greater than 0. To find the suitable coefficients we solve the ILP problem given by Eq. (8). The ILP could be solved with two substitution variables $s_{1}$ and $s_{2}$, giving the following QUBO encoding:
@@
To encode the multiplication inverse of GF($2^{4}$), we follow a similar procedure. In this case we are dealing with a map with a single four-bit input and a four-bit output: $(x_{1},x_{0},x_{2},x_{3})\rightarrow(z_{3},z_{2},z_{1},z_{0})$. Therefore, the bit-string $\mathbf{x}$ (again, $\mathbf{x}$ labels the input of the polynomial $g$) has 88 elements, for which we could solve the associated ILP problem with a single substitution variable $s$, giving a QUBO form:
@@
Table 3: The number of variables in the QUBO encoding of several Cryptography Functions. The results also indicate the sparseness of the resulting QUBO matrix (the density is calculated as a ratio of the nonzero elements in the QUBO matrix and $N^{2}$, with NN being the number of variables), and the magnitude of the largest element in the matrix. The encoding results for MD5, SHA1 and SHA-256 are presented for 22 cases: (a) using blocks of size $B=6$ for modular addition and (b) $B=1$ with k-SAT/ANF forms partitioned into at most $13$ clauses (to keep the coefficients low), see the main text for explanation. For encoding with $B=6$ and for AES there was no limitation on the length of the SAT forms.
@@ 4.2 Encoding constructions of MD5, SHA-1 and SHA-256 hash functions
As the main result of our work is the QUBO encoding of the AES we omit the detailed mathematical description of these hash algorithms. We just mention that they rely on a modular addition over GF($2^{32}$) as well as 33 and 44-input XOR operations, ternary conditional functions (ternary conditional function operates on 33 inputs: one binary input is used to choose between the two remaining inputs.) and majority conditional functions (MCF), like the one defined on 33 input variables: $MCF(x,y,z)=(x\wedge y)\vee(y\wedge z)\vee(z\wedge x)$. In other words, the MCF returns 11 if at least two of the input variables are 11, and 0 otherwise. The key point in encoding the mentioned hash functions is the efficient encoding of these components. The 44-input XOR (used in SHA-256) with substitution variable encoding its result (see Eq. (25)) can be encoded with the procedure described in Sec. $3.$ For the 33-input XOR formula $z\leftrightarrow x_{0}\oplus x_{1}\oplus x_{2}$ we provide an alternative QUBO encoding with a single substitution variable $(x_{0}+x_{1}+x_{2}-z+2s-2)^{2}$ obtained by ILP. The weak substitution variable takes on the value 11 and minimizes the formula when $x_{0}+x_{1}+x_{2}+1-z=1$. The ternary selection $z\leftrightarrow(x_{0}\wedge x_{1})\vee(\overline{x_{0}}\wedge x_{2})$ is encoded via a QUBO form:
@@
The encoding obtained by solving the associated ILP resulted in a single substitution variable $s$. It takes on the value 11 when $x_{0}+x_{1}+x_{2}+1-z=1$. Additionally, the 33-input majority function $z\leftrightarrow x_{0}\wedge x_{1}\vee x_{0}\wedge x_{2}\vee x_{1}\wedge x_{2}$ can be encoded without any substitution variable via the QUBO form
@@
The substitution variable $s$ takes on the value 11 and minimizes the formula when $\overline{x_{1}}\wedge(x_{0}+x_{2}+x_{3}\overset{?}{=}1)$ is True. Finally, only the discussion to encode the modular addition remains. Generally speaking, the most efficient way to encode addition in a QUBO is to write the addition as an equality $a+b=c$, then squaring the difference of the left and right sides and take the minimum of the quadratic function: $(a+b-c)^{2}\rightarrow 0$. When expanding $a$, $b$ and $c$ into binary formats, we obtain the QUBO representation of the addition. The drawback of this simple approach are the large coefficients in front of the most significant bit-variables. When dealing with $32$-bit addition relevant in the case of the studied hash functions, these coefficients become to large to be applicable on QUBO solving hardware. To address this issue, Ref. [12] developed a block-wise strategy to break the QUBO encoding of additions into smaller blocks. The QUBO formulae of the blocks are inter-connected via carry bits, as the sum of the terms might overflow the range covered by the bits of the block. Following this concept, we split the inputs into blocks by
@@
where $B\geq 1$ ($B\in\mathbb{N}$) is the block-size. With other words, the value of $B$ tells us the number of ”ordinary” bits encompassed via a single coefficient $a_{i}$ ($0\leq i\leq i_{max}-1$ with $i_{max}=\lceil\log_{2}(a)\rceil/B$, given that $B$ divides $\lceil\log_{2}(a)\rceil$ with no remainder) in a range $0\leq a_{i}\leq 2^{B}-1$. By expanding $b$ and $c$ as well, the addition reads as
@@
At this point, however, we need to account for potential overflows in the additions $a_{i}+b_{i}-c_{i}$ as none of the coefficients is allowed to get larger than $2^{B}-1$. This can be accomplished by introducing carry (or substitution) variables labeled by $s_{i}$, encoding the potential overflow of the sub-addition $a_{i}+b_{i}$. With this logic Eq. (46) is equivalent to the formulation
@@
In the MD5 algorithm the modular additions can be grouped into a single four-input and one two-input addition over 32-bit numbers in each round. The addition with 44 inputs implies the modification of (47) to
[…]
@@
describing the four-term addition $a+b+c+d=e$. The 44 terms in the addition imply 22 extra bits in each block to cover the potential overflow in the addition. The lower bound of maximal coefficient in the QUBO encoding due to expanding the squares lead to $2^{2B+2}$. We have also tried to further simplify the encoding logic with the ILP strategy developed in Sec. 2, though did not find any better encoding with less variables. In the actual MD5 implementation some pre-computed constants are also used in the procedure opening the door for possible reductions in the variable count by substitution constant bits and merging QUBO terms. Because of this, the coefficient in front of the merged QUBO terms are expected to further increase by some extent. We summarize our results on MD5 QUBO encoding in Table. 3. The encoding with $B=6$ (MD5a) seems to be more efficient than with $B=1$ (MD5b). This is because the $B=1$ case involves roughly 66-times more carry bits than the encoding with $B=6$ sized blocks. Also, a limit on the maximal length of ANF forms (i.e. the maximal number of inputs in XOR operations) has been set to $k=26$ for the case of $B=1$. (For the $B=6$ case and for the AES QUBO encodings we left the length of the SAT forms unlimited.) This little adjustment can be also used to decrease the coefficients in the QUBO instance, as the substitution variables would need to cover smaller ranges. This good practice becomes especially valuable when needed to port the QUBO instance onto a machine with limited numerical precision. Longer ANF forms are partitioned into smaller segments via substitution variables according to the procedure in Sec. $3.$ Finally, we need to mention that savings in the variable count usually comes with the price of larger coefficients, as reported in Table. 3. To run calculations on an actual hardware a reasonable trade-off needs to be found between QUBO variable count and the magnitude of the coefficients. Comparing to the QUBO encoding of Ref. [36] having $17280$ QUBO variables, our encoding strategy brings in a significant reduction to $10272$ QUBO variables obtained with $B=6$ sized blocks for addition.